Find an equation of the line having the given slope and containing the given point.
m=7/8, (1,-2)

Answer by DrBeeee(684)

General equation of a straight line is
y = mx + b. this is called the slope-intercept form because m stands for the slope and b is the y-intercept. This line passes through the point (0,b)
You want m=7/8, so you want
y = (7/8)x +b
To get the value of b, substitute the cordinates of the point (1,-2) giving
-2 = (7/8)(1) + b
Now solve for b
b = -2 - 7/8
b = -23/8
Your final equation is
y = (7/8)x - 23/8
Does this go through (1,-2)?
Does [-2 = (7/8)(1) - 23/8]?
Does [-2 = 7/8-23/8]?
Does [-2 = -16/8]?
Does [-2 = -2]? Yes
Answer: y = (7/8)x - 23/8